how to solve xy''+y'=0


1
Expert's answer
2021-06-28T16:41:33-0400
"xy''+y'=0"

"y'=u, y''=u'"

"xu'+u=0"

"x\dfrac{du}{dx}=-u"

"\dfrac{du}{u}=-\dfrac{dx}{x}"

"\int\dfrac{du}{u}=-\int\dfrac{dx}{x}"

"\ln|u|=-\ln|x|+\ln C_1"

"u=\dfrac{C_1}{x}"

"y'=\dfrac{C_1}{x}"

"\int dy=\int\dfrac{C_1}{x}dx"

"y=C_1\ln |x|+C_2"